OpenStudy (anonymous):

What polynomial identity should be used to prove that 19 = 27 - 8?

OpenStudy (anonymous):

@hero sorry again but this is my cousin and he needed help

hero (hero):

\(27 = 3^3\) right? \(8 = 2^3\) right? And \(27 - 8 = 3^3 - 2^3\) correct?

OpenStudy (anonymous):

yes

hero (hero):

So what kind of polynomial identity could we use for this?

OpenStudy (anonymous):

difference of cubes?

hero (hero):

Correct. Great job @xiphoidx

OpenStudy (anonymous):

Omg thanks so much! I was always confused with how to do this @hero